article with acharge of 1.0 pC d a masof3.0 ×106kg is released from rest at point A and accelerates toward point Be among there with a speed of 39m/s. The only force ating on the particle is the ele tre foot (a) Which point is at the higher potential ● point B Give your reasoning Negative charge accelerates from a lower potential to a higher potential Negative charge accelerates from a higher potential to a lower potential (b) What is the potentialdifference Ve- Va between A and B dditional Materials Section 192
Answer #1

answer ) a) the correct option is point B

reasoning -negative charge particles accelerates from a lower potential to higher potential.

b) here the mechanical energy is conserved which means

\Deltak+\Deltau=0

\Deltak=-\Deltau

\Deltak=1/2mv b2-1/2mva2=1/2mvb2( since va=0)

so \Deltak=-\Deltau=-q\Deltav

\Deltav=-mvb2/2q=-3*10-6*392/2*-1.6*10-6=1426V

so the answer is 1426V